OpenStudy (anonymous):

A Country's population in 1991 was 15 million. In 2001 it was 16 million. Estimate the population in 2005 using the exponential growth formula. Round your answer to the nearest million

OpenStudy (anonymous):

a=16,000,000−15,000,0002001−19912001−1991=1,000,00010=100,000 y=100,000∗bx Thus[16,000,000=100,000*b^{10}\] b10=160